OpenStudy (anonymous):

13.4 grams of iron (III) oxide reacts with 12.3 grams of carbon monoxide to produce 8.23 g of pure iron, what are the theoretical yield and percent yield of this reaction? Be sure to show the work that you did to solve this problem.Please Help

OpenStudy (aaronq):

1. write a reaction then balance it. 2. find the moles of iron produced by using the stoichiometric coefficients. e.g. for a general reaction: aA + bB \(\rightleftharpoons\) cC where upper case are the species, and lower case are the coefficients, \(\dfrac{n_A}{a}=\dfrac{n_B}{b}=\dfrac{n_C}{c}\) from here, you can solve for moles then convert to grams. The amount of grams you find here is the theoretical yield. 3. finding percent yield, use: \(\% yield=\dfrac{actual\;yield}{theoretical\;yield}*100\%\) the units should both be the same, otherwise it won't work.

OpenStudy (anonymous):

Fe2O3 + 3 CO-->2 Fe + 3 CO2

OpenStudy (anonymous):

right?

OpenStudy (aaronq):

yup, you now need to convert the masses given to moles and use their stoichiometric coefficients to build a ratio, like i stated above.

OpenStudy (anonymous):

can u help with that please

OpenStudy (aaronq):

convert to moles first

OpenStudy (anonymous):

idk how

OpenStudy (aaronq):

13.4 grams of iron (III) oxide reacts with 12.3 grams of carbon monoxide \(n=\dfrac{m}{M}\) where, n=moles, m=mass (in g), M=molar mass (in g/mol). You'll find M on the periodic table

OpenStudy (anonymous):

...

OpenStudy (aaronq):

...

OpenStudy (aaronq):

?

OpenStudy (anonymous):

can you just do the question out so i can find the percent yield

OpenStudy (aaronq):

i'm not going to do your work

OpenStudy (anonymous):

I just need explanation, like a bit of a back and forth thing.

OpenStudy (aaronq):

i wrote above how you can find the amount of moles of both reactants.

OpenStudy (anonymous):

or can u add someone else to help me please?

OpenStudy (anonymous):

ill give a medal

OpenStudy (aaronq):

i'm trying to help you but you don't wanna do any work.

OpenStudy (anonymous):

@Hero please help

OpenStudy (anonymous):

@Chemistry please help